# Robot Navigation in Unknown and Cluttered Workspace with Dynamical System Modulation in Starshaped Roadmap

### v2

@@ Abstract
Compared to conventional decomposition methods that use ellipses or polygons to represent free space, starshaped representation can better capture the natural distribution of sensor data, thereby exploiting a larger portion of traversable space. This paper introduces a novel motion planning and control framework for navigating robots in unknown and cluttered environments using a dynamically constructed starshaped roadmap. Our approach generates a starshaped representation of the surrounding free space from real-time sensor data using piece-wise polynomials. Additionally, an incremental roadmap maintaining the connectivity information is constructed, and a searching algorithm efficiently selects short-term goals on this roadmap. Importantly, this framework addresses deadend situations with a graph updating mechanism. To ensure safe and efficient movement within the starshaped roadmap, we propose a reactive controller based on Dynamic System Modulation (DSM). This controller facilitates smooth motion within starshaped regions and their intersections, avoiding conservative and short-sighted behaviors and allowing the system to handle intricate obstacle configurations in unknown and cluttered environments. Comprehensive evaluations in both simulations and real-world experiments show that the proposed method achieves higher success rates and reduced travel times compared to other methods. It effectively manages intricate obstacle configurations, avoiding conservative and myopic behaviors. The source code will be released on website [^1] Available at: github.com/kkkkkaiai/starshaped_roadmap.
@@ I Introduction
Figure 1: Overview of the proposed navigation framework. (a) In our method, a dynamic starshaped roadmap is constructed in real-time to facilitate the proposed reactive controller based on DSM, enabling the robot to navigate smoothly towards its goal. (b) Using our proposed method, the robot successfully navigates through unknown and cluttered environments, demonstrating its effectiveness in real-world scenarios. (c)-(e) Comparing with the representation of free regions by ellipses and convex polygons, our approach recognizes them as starshaped regions, which capture a larger area of the free space, so that the perception information can be fully exploited for robot navigation.
@@
  • We propose a novel method for representing traversable space in unknown and cluttered environments by constructing starshaped regions from locally perceived point cloud data. This approach maximizes the use of available free space, avoiding conservative behavior, and dynamically generates a roadmap with extendable regions.
@@ III-B Starshaped Dynamic Roadmap Generation
In this work, our roadmap serves a dual purpose: it identifies potential paths for robot navigation and delineates the traversable areas within the space. The proposed roadmap, called starshaped roadmap, consists of nodes and edges. Each node stores a position and the traversable area generated from that position as the origin, representing the range within which the robot can move without collision. The edges define the direct connections between nodes, representing feasible transitions within the environment.
@@ III-B1 Representation of the Traversable Regions
Subsequently, we employ piece-wise polynomials to describe the boundary. We first transform the point cloud $\mathcal{L}_{\mathbf{p}_{ref}}$ from the Cartesian coordinate system into the polar coordinate system, with the origin of the point cloud at $\mathbf{p}_{ref}$. One key advantage of using the polar coordinate system is that any fitting inaccuracies only affect the precision of the boundary and not alter its compliance with the starshaped definition. Each point $\mathbf{p}_{l}=(p_{x,l},p_{y,l})\in\mathcal{L}$ to the polar coordinates $(\theta,d)$, where $\theta$ represents the angular direction and and $d$ denotes the radial distance from $\mathbf{p}_{ref}$. For any position within the region, excluding $\mathbf{p}_{ref}$, the transformation is given by the following formulation:
@@ III-C3 Motion Control in Starshaped Roadmap
where $\mathbf{r}(\mathbf{p})=\mathbf{p}_{ref}-\mathbf{p}$ is the eigenvector pointing toward to the reference point, associated with the scaling factor $\lambda_{r}(\mathbf{p})$, which governs the reduction of velocity in the critical direction towards the boundary. The remaining eigenvectors $\mathbf{e}_{1}(\mathbf{p}),\dots,\mathbf{e}_{n-1}(\mathbf{p})$ are orthogonal to $\mathbf{r}(\mathbf{p})$ and are associated with the scaling factor $\lambda_{e}(\mathbf{p})$, which allows greater freedom of motion in safer directions.
@@
Figure 3: Ablation study of the proposed adaptive segmentation method for piecewise polynomials. Compared to different fixed segment numbers, our method achieves higher accuracy while offering a trade-off in computational time.
@@ IV Experiments and Results
To improve the accuracy of boundary fitting for starshaped regions in Section III-B1, we utilize the rate of change in radial distance to adaptively determine the polynomial segments, balancing accuracy and computation time. In this case, $N=3$. The results, compared to those obtained using a fixed number of segments, are shown in Fig. 3.
@@ IV-A2 Maze Scenario
In terms of computation time, the average result for controlling the robot is $3.98\pm 0.96$ ms, and $4.37\pm 1.16$ ms in the forest and maze scenarios, respectively. The computational efficiency allows us to set a high control frequency, more than 200 Hz, for the robot to perform rapid reactions. The fitting process for the starshaped representation takes approximately $8\pm 1.3$ ms, as shown in Fig. 3, and it only needs to be executed when expanding the roadmap and therefore has minimal impact on the overall computational efficiency of the proposed algorithm.
@@ IV-B Real-World Experiment
To validate our method in real-world conditions, we conduct experiments using a differential drive mobile robot equipped with a 2D LiDAR for environmental perception. The robot navigates through a $6\,\text{m}\times 8\,\text{m}$ area populated with various obstacles. In the experiment, depicted in Fig. 6, the robot successfully maneuvers around randomly placed obstacles, reaching its destination. It takes an average of 10 ms to construct starshaped regions from the sensor data incrementally. The robot maintains an average speed of 0.4 m/s, with control inputs calculated in 3 ms on average. Between 8 and 12 s, as shown in Fig. 6(b), the robot gets stuck. The roadmap is updated immediately and identifies a new short-term goal to bypass the obstacle.

### v3

@@ Abstract
Compared to conventional decomposition methods that use ellipses or polygons to represent free space, starshaped representation can better capture the natural distribution of sensor data, thereby exploiting a larger portion of traversable space. This paper introduces a novel motion planning and control framework for navigating robots in unknown and cluttered environments using a dynamically constructed starshaped roadmap. Our approach generates a starshaped representation of the surrounding free space from real-time sensor data using piece-wise polynomials. Additionally, an incremental roadmap maintaining the connectivity information is constructed, and a searching algorithm efficiently selects short-term goals on this roadmap. Importantly, this framework addresses dead-end situations with a graph updating mechanism. To ensure safe and efficient movement within the starshaped roadmap, we propose a reactive controller based on Dynamic System Modulation (DSM). This controller facilitates smooth motion within starshaped regions and their intersections, avoiding conservative and short-sighted behaviors and allowing the system to handle intricate obstacle configurations in unknown and cluttered environments. Comprehensive evaluations in both simulations and real-world experiments show that the proposed method achieves higher success rates and reduced travel times compared to other methods. It effectively manages intricate obstacle configurations, avoiding conservative and myopic behaviors. The source code will be released on website [^1] Available at: github.com/kkkkkaiai/starshaped_roadmap.
@@ I Introduction
Figure 1: Highlight of the proposed navigation framework. (a) Our method, a dynamic starshaped roadmap is constructed in real-time to facilitate the proposed reactive controller based on DSM, enabling the robot to navigate smoothly towards its goal. (b) With our approach, the robot successfully navigates through unknown and cluttered environments, demonstrating its effectiveness in real-world scenarios. (c)-(e) Comparing with the representation of free regions by ellipses and convex polygons, our approach uses starshaped regions, covering a larger area and enabling better use of perception information for navigation.
@@
  • We propose a novel method for representing traversable space in unknown and cluttered environments by constructing starshaped regions from locally perceived point cloud data. This approach maximizes the use of available free space, avoids conservative behavior, and dynamically generates a roadmap with extendable regions.
@@ III-B Starshaped Dynamic Roadmap Generation
In this work, our roadmap serves two key purposes: it identifies potential paths for robot navigation and delineates the traversable areas within the space. The proposed roadmap, called starshaped roadmap, consists of nodes and edges. Each node stores a position and the traversable area generated from that position as the origin, representing the range within which the robot can move without collision. The edges define the direct connections between nodes, representing feasible transitions within the environment.
@@ III-B1 Representation of the Traversable Regions
Subsequently, we employ piece-wise polynomials to describe the boundary. We first transform the point cloud $\mathcal{L}_{\mathbf{p}_{ref}}$ from the Cartesian coordinate system into the polar coordinate system, with the origin of the point cloud at $\mathbf{p}_{ref}$. One key advantage of using the polar coordinate system is that any fitting inaccuracies only affect the precision of the boundary and do not alter its compliance with the starshaped definition. Each point $\mathbf{p}_{l}=(p_{x,l},p_{y,l})\in\mathcal{L}$ to the polar coordinates $(\theta,d)$, where $\theta$ represents the angular direction and and $d$ denotes the radial distance from $\mathbf{p}_{ref}$. For any position within the region, excluding $\mathbf{p}_{ref}$, the transformation is given by the following formulation:
@@ III-C3 Motion Control in Starshaped Roadmap
where $\mathbf{r}(\mathbf{p})=\mathbf{p}_{ref}-\mathbf{p}$ is the eigenvector pointing toward the reference point, associated with the scaling factor $\lambda_{r}(\mathbf{p})$, which governs the reduction of velocity in the critical direction towards the boundary. The remaining eigenvectors $\mathbf{e}_{1}(\mathbf{p}),\dots,\mathbf{e}_{n-1}(\mathbf{p})$ are orthogonal to $\mathbf{r}(\mathbf{p})$ and are associated with the scaling factor $\lambda_{e}(\mathbf{p})$, which allows greater freedom of motion in safer directions.
@@
Figure 3: Ablation study of the proposed adaptive segmentation method for piecewise polynomials. Compared to different fixed segment numbers, our method achieves higher accuracy while offering a trade-off in computation time.
@@ IV Experiments and Results
To improve the accuracy of boundary fitting for starshaped regions in Section III-B1, we utilize the rate of change in radial distance to adaptively determine the polynomial segments, balancing accuracy and computation time. In this case, $N=3$. The error performance, in comparison with those obtained using a fixed number of segments, are presented in Fig. 3.
@@ IV-A2 Maze Scenario
IV-A3 Computation Time
The average computing time for controlling the robot is $3.98\pm 0.96$ ms, and $4.37\pm 1.16$ ms in the forest and maze scenarios, respectively. The computational efficiency allows us to set a high control frequency, more than 200 Hz, for the robot to perform rapid reactions. The fitting process for the starshaped representation takes approximately $8\pm 1.3$ ms, as shown in Fig. 3, and it only needs to be executed when expanding the roadmap and therefore has minimal impact on the overall computational efficiency of the proposed algorithm.
@@ IV-B Real-World Experiment
To validate our method in real-world conditions, we conduct experiments using a differential drive mobile robot equipped with a 2D LiDAR for environmental perception. The robot navigates through a $6\,\text{m}\times 8\,\text{m}$ area populated with various obstacles. In the experiment, depicted in Fig. 6, the robot successfully maneuvers around randomly placed obstacles, reaching its destination. It takes an average of 10 ms to construct starshaped regions from the sensor data incrementally. The robot maintains an average speed of 0.4 m/s, with control inputs calculated in 3 ms on average. The robot gets stuck around 12 s, at which point the roadmap is updated immediately to identify a new short-term goal, allowing the robot to bypass the obstacle.